# Algebraic non-Hermitian skin effect and unified non-Bloch band theory in arbitrary dimensions

### v1

@@ Abstract
The non-Hermitian skin effect, characterized by a proliferation of exponentially-localized edge modes in open-boundary systems, has led to the discovery of numerous novel physical phenomena that challenge the limits of conventional band theory. In sharp contrast to the traditional exponential localization, this manuscript reports a new kind of non-Hermitian skin effect, which we term the “algebraic non-Hermitian skin effect.” This effect emerges across a diverse spectrum of non-Hermitian systems in both two- and higher space dimensions. For 2D systems with algebraic non-Hermitian skin effect, on geometries such as a torus or cylinder, these systems exhibit behavior reminiscent of the conventional non-Hermitian skin effect, where eigenmodes are either bulk Bloch waves (on a torus) or exponentially localized edge modes (on a cylinder). However, if the same system is placed on a disk or any geometrical shape featuring open boundaries in all directions, the skin modes immediately transform into the algebraic form, with amplitude decaying as a power-law function of the distance from the boundary. This power-law decay, coupled with the sensitivity of edge mode decay to the real-space topology, cannot be characterized within the existing theoretical framework of the generalized Brillouin zone (GBZ). To explore these novel effects, we formulate a unified GBZ framework that is universally applicable to all variations of non-Hermitian skin effects across any spatial dimension, developed through the usage of a generalized transfer-matrix approach. Within this new framework, we find that in a $d$-dimensional non-Hermitian system, in general, the GBZ manifold’s dimensionality must fall into the range from $d$ to $12d-$ 1, denoted by ${d\leq\dim\text{GBZ}\leq 2d-1}$. In 1D, this inequality is trivial because the upper and lower bounds converge, forcing the GBZ’s dimensionality to match with that of the physical space, in agreement with existing knowledge on the exponential non-Hermitian skin effect. However, in 2D and above, this inequality indicates that there is no obligation for the GBZ’s dimensionality to concur with the physical space’s dimensionality, which gives rise to a new class of non-Hermitian skin effects, specifically, the algebraic non-Hermitian skin effect. We have implemented this framework on tight-binding Hamiltonians to exemplify our findings within both two- and three-dimensional settings. This investigation not only unveils a novel category of the non-Hermitian skin effect but also offers a comprehensive theoretical structure that describes skin effects in any non-Hermitian system, irrespective of spatial dimensionality.
@@ I Introduction
For systems that interact with external environments, the utilization of non-Hermitian Hamiltonians has proven effective in encapsulating core physical properties of these systems Ashida et al. (2020); Bergholtz et al. (2021); Ding et al. (2022). These non-Hermitian frameworks have been increasingly recognized for their utility in open classical-wave systems —including photonic Zhen et al. (2015); Zhou et al. (2018); Cerjan et al. (2019), acoustic Huang et al. (2024); Zhang et al. (2021a), and mechanical metamaterials Scheibner et al. (2020a, b); Zhou and Zhang (2020); Banerjee et al. (2021); Shankar et al. (2022)— as well as in open quantum systems Song et al. (2019); Liu et al. (2020); Haga et al. (2021); Xue et al. (2022). In recent years, research interest in non-Hermitian systems has surged as a result of their unique and distinct properties: non-Hermitian Hamiltonians allow complex-valued eigenvalues and relieve the requirement for eigenstates to be orthogonal to each other Brody (2013). This deviation from the conventional norms enriched our understanding and led to the emergence of new insights and principles in physics beyond the scope of traditional band theory Hatano and Nelson (1996); Esaki et al. (2011); Shen et al. (2018); Gong et al. (2018); Nagai et al. (2020).
@@
One striking phenomenon of non-Hermitian band systems is the non-Hermitian skin effect Yao and Wang (2018); Kunst et al. (2018); Yokomizo and Murakami (2019); Lee and Thomale (2019); Borgnia et al. (2020); Zhang et al. (2020); Okuma et al. (2020); Yang et al. (2020); Kawabata et al. (2020a); Longhi (2019); Deng and Yi (2019); Xiao et al. (2020); Helbig et al. (2020); Ghatak et al. (2020); Li et al. (2020a); Claes and Hughes (2021); Lu et al. (2021); Li et al. (2023); Gu et al. (2022); Cheng et al. (2024); Ochkan et al. (2024); Okuma and Sato (2023); Lin et al. (2023); Zhang et al. (2022a), which is characterized by the localization of almost all eigenstates at the open boundaries Lee (2016); Martinez Alvarez et al. (2018), and a notable distinction between spectra under periodic and open boundary conditions Xiong (2018). This effect defies traditional band theory and compels a reevaluation of well-established principles, such as the bulk-boundary correspondence that characterizes Hermitian systems Chiu et al. (2016); Kawabata et al. (2019); Xiao et al. (2020); Helbig et al. (2020); Ghatak et al. (2020). Furthermore, the non-Hermitian skin effect spawns a variety of fascinating phenomena unique to non-Hermitian systems, including non-reciprocal responses Yi and Yang (2020); Xue et al. (2021); Li et al. (2021), directional wavepacket invisibility Longhi et al. (2015); McDonald et al. (2018); Lee et al. (2019a); Bessho and Sato (2021); Liang et al. (2022); Zhang et al. (2023), and improved sensitivity in sensors that scales with the size of the system Budich and Bergholtz (2020); McDonald and Clerk (2020); Bao et al. (2022), among others Longhi (2022); Gong et al. (2022); Fang et al. (2023). Such advancements hold promise for robust and adaptable control over wave and signal transmission Weidemann et al. (2020); Wang et al. (2022a). To fully comprehend the non-Hermitian skin effect, it is necessary to consider an extension of Bloch wavevectors into the complex plane, thereby advancing Bloch band theory into the more encompassing non-Bloch band paradigm Yao and Wang (2018); Yokomizo and Murakami (2019); Yang et al. (2020); Song et al. (2019); Deng and Yi (2019); Kawabata et al. (2020b); Helbig et al. (2020); Ghatak et al. (2020); Xiao et al. (2020). This evolution of the theory accounts for the emergent complexities and novel behaviors witnessed in non-Hermitian systems.
@@
The non-Hermitian skin effect in one dimension has been well-established through the GBZ framework. By mapping the Bloch wavevector $k$ to $\beta=e^{ik}$, the Brillouin zone (BZ) is mapped into a unit circle in the complex-$\beta$ plane as $k$ varies from $-\pi$ to $\pi$. To characterize the non-Hermitian skin effect, the wavevectors $k$ are extended to complex values, and accordingly, the BZ is generalized to the GBZ, represented as piecewise analytic closed loops of $\beta$ in the complex-$\beta$ plane Yao and Wang (2018); Yokomizo and Murakami (2019); Zhang et al. (2020); Yang et al. (2020). The GBZ accurately captures the localization of the skin effect, with the radius of the GBZ indicating its localization length and direction. Within this framework, the 1D non-Hermitian skin modes can be interpreted as Bloch waves modified by exponential localization prefactors. A local gauge transformation can be applied to transform these skin modes back to conventional Bloch waves, by removing the exponential prefactors Hatano and Nelson (1996); Yao and Wang (2018); Lee and Thomale (2019). This illustrates that the 1D non-Bloch band paradigm is not far from the traditional Bloch band framework, facilitated by the exponential localization law of the 1D non-Hermitian skin effect.
@@
However, in two and higher dimensions, the exponential localization characteristic of the non-Hermitian skin effect needs reevaluation, due to the lack of an accurate high-dimensional GBZ description and the following observations. A direct extension of the 1D skin effect would be that the higher-dimensional skin effects also exhibit the same exponential localization behavior; accordingly, these skin modes should localize at the corners of open boundaries under the thermodynamic limit Yao et al. (2018); Lee et al. (2019b); Liu et al. (2019); Li et al. (2020b); Zou et al. (2021); Zhang et al. (2021b). However, recent numerical and experimental observations have indicated that the skin modes universally accumulate on edges rather than corners of open-boundary geometries Zhang et al. (2022b); Wang et al. (2022b); Fang et al. (2022); Zhou et al. (2023); Wang et al. (2023a); Wan et al. (2023); Qin et al. (2024); Zhang et al. (2024); Wang et al. (2023b); Zhao et al. (2023), representing a phenomenon without a direct 1D counterpart and lacking a unified GBZ description. Additionally, in higher dimensions, the complex open-boundary eigenenergies can cover a region of finite areas in the complex energy plane, indicating a need for a fundamentally different GBZ formulation compared to 1D systems. Given these considerations, we pose the following key questions: (i) Does the higher-dimensional skin effect inherently exhibit exponential localization? If not, what are the governing principles of its localization? (ii) Is it possible to develop a unified GBZ framework that accurately describes all forms of non-Hermitian skin effects? Although attempts have been made to adapt the GBZ formula to higher dimensions Yao et al. (2018); Yokomizo and Murakami (2023); Jiang and Lee (2023); Wang et al. (2024); Hu (2023); Zhang et al. (2024); Xu et al. (2023), achieving a unified GBZ framework that incorporates all critical aspects and precisely represents all variants of non-Hermitian skin effects continues to be a challenge and is still under debate.
@@
In this manuscript, we report the algebraic non-Hermitian skin effect in two and higher dimensions. Under open boundary conditions (OBCs), the amplitudes of skin modes in these systems exhibit power-law decay, contrasting sharply with the exponential decay commonly seen in 1D systems. However, when the system’s geometry is modified into a cylindrical shape, the decay of the skin modes transitions to exponential. This power-law decay and the sensitivity of the decay pattern to boundary conditions are highly unconventional. In Hermitian systems, comparable behavior is only observed for very special systems under specific and fine-tuned conditions: for instance, in 2D elasticity with zero bulk modulus Sun et al. (2012). In these Hermitian systems, power-law and exponential edge modes arise due to a special characteristic of the underlying system—an emergent conformal symmetry which supports indefinitely many conserved quantities. Given that 2D conformal mappings are highly sensitive to the topology of the underlying geometry, they naturally give rise to exponential functions for a cylindrical geometry and power-law functions for a disk or annulus geometry Sun et al. (2012). Any deviation from the conformal limit eliminates this effect in these Hermitian systems. In contrast, in non-Hermitian systems, similar phenomena occur from a totally different origin, without the need for specialized conditions or fine-tuning, and there is no requirement for conformal symmetry to observe such phenomena. The presence of generic and robust algebraic edge modes represents a unique feature of non-Hermitian systems in 2D and above, and stands in stark contrast to what can be achieved in Hermitian setups and/or 1D non-Hermitian systems.
@@
This finding of robust algebraic non-Hermitian skin effect is supported by establishing a unified GBZ framework that applies to any non-Hermitian system, regardless of its spatial dimensionality. The unified GBZ framework is established based on a generalized transfer-matrix approach. We demonstrate that the GBZ’s dimensionality depends on the coverage of the open-boundary continuum spectrum, and the number of GBZ constraints. In high dimensions, the GBZ constraint can be fewer than the space dimensionality, meaning that standing waves can be formed in either one direction or multiple directions, without necessarily forming along all space dimensions. According to the case of the open-boundary spectrum and the number of standing-wave conditions, we find that in $d$-dimensional lattice systems, the dimension of GBZ manifold ranges from $d$ to ${2d-1}$. We have also implemented this framework on tight-binding Hamiltonians in two and three dimensions, successfully demonstrating these new phenomena. The outcomes are in complete accord with numerical simulations, exhibiting consistency both quantitatively and qualitatively.
@@
The remainder of the paper is organized as follows. In Sec. II, we summarize key concepts, focusing on dimensional counting to explain the emergence of algebraic skin modes in $d>1$ systems and their transition to exponential decay in cylindrical geometries, without delving into detailed theory analysis. In Sec. III, we introduce rigorous methodologies to derive the unified GBZ formula by employing the non-Bloch transfer matrix approach. With this exact theoretical method, we construct the GBZ, compute the open boundary spectrum/eigenstates, and prove the dimension analysis and theoretical framework outlined in Sec. II. In Sec. IV, utilizing this new GBZ formula, we prove that the reciprocal non-Hermitian skin effect is the algebraic skin effect, where the amplitudes of skin modes follow a power-law decay, instead of the conventionally expected exponential law. In Sec. V, we generalize the GBZ formulation into arbitrary $d$ dimensions, and provide examples in three dimensions.
@@ II Dimensionality analysis for the generalized Brillouin zone and summary of key concepts
Before delving into the universal GBZ formula in arbitrary spatial dimensions, we commence with an analysis of dimensionality regarding BZ and GBZ. In this section, our discussion pivots solely on qualitative aspects, leaving the meticulous formulation of the GBZ theory to be explored in the subsequent section.
@@
In the construction of eigenstates for a given eigenenergy $E$, one needs to consider only a specific subset of the BZ in Hermitian setups or GBZ in non-Hermitian systems. This relevant subset will be termed the Fermi surface (FS) in the case of the BZ or the generalized Fermi surface (GFS) in connection with the GBZ. The dimensionality of the FS or GFS can be determined through the following relation:
@@
$\dim{\text{GFS}}$ is a key factor in determining the characteristics of skin modes. When $\dim\text{GFS}=0$, skin modes must decay exponentially with the increase in distance from the open boundaries. On the other hand, if $\dim\text{GFS}>0$, the decay of edge modes may follow different functional forms, such as power-law. Using the dimension counting provided above, it becomes evident that $\dim\text{GFS}>0$ can only occur in dimensions greater than one, $d>1$. Consequently, this implies that algebraic non-Hermitian skin effects cannot manifest in one-dimensional systems, $d=1$.

### v2

# Algebraic non-Hermitian skin effect and generalized Fermi surface formula in arbitrary dimensions

@@ Abstract
The non-Hermitian skin effect, characterized by a proliferation of exponentially localized edge modes in open-boundary systems, has led to the discovery of numerous novel physical phenomena that challenge the limits of conventional band theory. In sharp contrast to this familiar exponential localization, we report a distinct phenomenon—the algebraic non-Hermitian skin effect—which arises generically in non-Hermitian systems with two or more spatial dimensions. In such cases, the amplitude of skin modes typically decays from the boundary following a power law, rather than an exponential form—a behavior not captured by existing theoretical frameworks. To bridge this gap and describe the transition in localization from one to higher dimensions, we develop a generalized Fermi surface framework applicable to open-boundary systems in arbitrary dimensions. This framework not only reproduces known results for the exponential skin effect in 1D, but also predicts a new class of skin effects with algebraic decay in 2D and above. We demonstrate this framework in both tight-binding and continuum models in two and three dimensions. This investigation not only unveils a novel category of the non-Hermitian skin effect but also offers a comprehensive theoretical structure that describes skin effects in any non-Hermitian system, irrespective of its spatial dimensionality.
@@ I Introduction
For systems that interact with external environments, the utilization of non-Hermitian Hamiltonians has proven effective in encapsulating core physical properties of these systems [1, 2, 3]. These non-Hermitian frameworks have been increasingly recognized for their utility in open classical-wave systems —including photonic [4, 5, 6], acoustic [7, 8], and mechanical metamaterials [9, 10, 11, 12, 13]— as well as in open quantum systems [14, 15, 16, 17, 18]. In recent years, research interest in non-Hermitian systems has surged as a result of their unique and distinct properties: non-Hermitian Hamiltonians allow complex-valued eigenvalues and relieve the requirement for eigenstates to be orthogonal to each other [19]. This deviation from the conventional norms enriched our understanding and led to the emergence of new insights and principles in physics beyond the scope of traditional band theory [20, 21, 22, 23, 24].
@@
One striking phenomenon of non-Hermitian band systems is the non-Hermitian skin effect [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51], which is characterized by the localization of almost all eigenstates at the open boundaries [52, 53], and a notable distinction between spectra under periodic and open boundary conditions (OBCs) [54]. This effect defies traditional band theory and compels a reevaluation of well-established principles, such as the bulk-boundary correspondence that characterizes Hermitian systems [55, 56, 36, 37, 38]. Furthermore, the non-Hermitian skin effect spawns a variety of fascinating phenomena unique to non-Hermitian systems, including non-reciprocal responses [57, 58, 59], directional wavepacket invisibility [60, 61, 62, 63, 64, 65], and improved sensitivity in sensors that scales with the size of the system [66, 67, 68], among others [69, 70, 71]. Such advancements hold promise for robust and adaptable control over wave and signal transmission [72, 73]. To fully comprehend the non-Hermitian skin effect, it is necessary to consider an extension of Bloch wavevectors into the complex plane, thereby advancing Bloch band theory into the more encompassing non-Bloch band paradigm [25, 27, 32, 14, 35, 74, 37, 38, 36]. This evolution of the theory accounts for the emergent complexities and novel behaviors witnessed in non-Hermitian systems.
@@
The non-Hermitian skin effect in one dimension has been well-established through the generalized Brillouin zone (GBZ) framework. By mapping the Bloch wavevector $k$ to $\beta=e^{ik}$, the Brillouin zone (BZ) is represented as the unit circle in the complex-$\beta$ plane as $k$ varies from $\pi$ to $\pi$. To characterize the non-Hermitian skin effect, the wavevectors $k$ are extended to complex values, and accordingly, the BZ is generalized to the GBZ, represented as piecewise analytic closed loops of $\beta$ in the complex-$\beta$ plane [25, 27, 30, 32]. The GBZ accurately captures the localization of the skin effect, with its radius indicating the localization length and direction. Within this framework, the 1D non-Hermitian skin modes can be interpreted as Bloch waves modulated by an exponentially localized envelope. A local imaginary gauge transformation can be applied to transform these skin modes back into conventional Bloch waves by removing the exponential prefactors [20, 25, 28]. This illustrates that the 1D non-Bloch band paradigm is not far from the traditional Bloch band framework, owing to the exponential localization law of the 1D non-Hermitian skin effect.
@@
However, in two and higher dimensions, the exponential localization characteristic of the non-Hermitian skin effect warrants reevaluation, due to the absence of an accurate high-dimensional GBZ description for wavefunctions and the following considerations. A straightforward extension of the 1D skin effect to higher dimensions would suggest that skin modes also exhibit exponential localization. As a result, these modes would accumulate at the corners of open-boundary systems in the thermodynamic limit [75, 76, 77, 78, 79, 80]. However, recent numerical and experimental observations have indicated that skin modes universally accumulate on edges rather than at corners of open-boundary geometries [81, 82, 83, 84, 85, 86, 87, 88, 89, 90], representing a phenomenon without a direct 1D counterpart and lacking a comprehensive theoretical framework for accurately describing higher-dimensional OBC wavefunctions. Additionally, the GBZ framework generally implies the formation of standing waves under OBCs. For instance, the 1D GBZ framework suggests that two non-Bloch waves with the same localization length can combine to form standing waves under OBCs, as illustrated by the 1D case in Fig. 1. However, extending this concept of standing wave formation to higher dimensions remains challenging [91] and is not yet fully understood, posing a significant barrier to understand and develop a comprehensive non-Bloch framework in higher dimensions. Given these considerations, we pose the following key questions: (i) Are high-dimensional skin effects generally characterized by short-range exponential localization? If not, what forms of localization behavior can they exhibit? (ii) Is it possible to develop a systematic theoretical framework that accurately captures higher-dimensional non-Hermitian skin effects in the thermodynamic limit? Although several attempts have been made to extend the GBZ formalism to higher dimensions [75, 91, 92, 93, 94, 88, 95], establishing a theoretical framework that can account for diverse localization behaviors, such as edge-localized skin effects, and accurately construct OBC wavefunctions remains a significant challenge and is still under active debate.
@@
In this paper, we report the algebraic non-Hermitian skin effect in two and higher dimensions. In contrast to the exponential localization of skin modes commonly observed in 1D systems, we find that the amplitudes of skin modes in higher dimensions exhibit power-law localization under OBCs. When the system’s geometry is modified into a cylindrical shape, the decay of the skin modes transitions to exponential. This power-law decay and the sensitivity of the decay pattern to boundary conditions are highly unconventional. In Hermitian systems, comparable behavior is only observed for very special systems under specific and fine-tuned conditions: for instance, in 2D elasticity with zero bulk modulus [96]. In these Hermitian systems, power-law and exponential edge modes arise due to a special characteristic of the underlying system—an emergent conformal symmetry which supports indefinitely many conserved quantities. Given that 2D conformal mappings are highly sensitive to the topology of the underlying geometry, they naturally give rise to exponential functions for a cylindrical geometry and power-law functions for a disk or annulus geometry [96]. Any deviation from the conformal limit eliminates this effect in these Hermitian systems. In contrast, in non-Hermitian systems, similar phenomena occur from a totally different origin, without the need for specialized conditions or fine-tuning, and there is no requirement for conformal symmetry to observe such phenomena. The presence of generic and robust algebraic edge modes represents a unique feature of non-Hermitian systems in 2D and above, and stands in stark contrast to what can be achieved in Hermitian setups and/or 1D non-Hermitian systems.
Figure 1:  Illustration of the standing wave formation in one, two, and $d$ dimensions. Here, $n$ denotes the number of enforced standing-wave conditions, with red double arrows illustrating the standing-wave directions and the blue wide arrows indicating the transfer directions. In one dimension, standing wave condition can only be enforced along the single spatial direction. In contrast, in two and higher dimensions, standing waves can form along one, multiple, or all spatial directions. In the case of $n=d$ (where $d$ is the spatial dimension), standing wave conditions fully constrain the system, a bulk formula independent of boundary conditions can be established. However, in more general cases where $n<d$, such as $n=1$ in two dimensions, the boundary conditions become significant.
@@
This finding of a robust algebraic non-Hermitian skin effect is supported by the establishment of a generalized Fermi surface (GFS) framework, which applies to generic non-Hermitian systems with regular open-boundary geometries. While the OBC in one dimension is unambiguous, corresponding to the ends of an open chain, the notion of OBC geometry becomes more diverse in higher dimensions. Here, regular boundary geometries specifically refer to $d$-dimensional parallelotopes, such as parallelograms in 2D and parallelepipeds in 3D. The GFS framework is established based on the standing wave formation, as illustrated in Fig. 1. In one dimension, the standing waves can form along the only one spatial dimension and make the $\dim\text{GBZ}={d}=1$ under the existing non-Bloch framework [25, 27]. While in two or higher $d$-dimensional systems, there are $d$ distinct scenarios in terms of the number of standing wave conditions as shown in Fig. 1. In general higher-dimensional cases, the GBZ constraint can be fewer than the space dimensionality, meaning that standing waves can be formed in either one direction or multiple directions, without necessarily forming along all space dimensions. We find that in $d$-dimensional lattice systems, the dimension of GBZ manifold ranges from $d$ to 2d1{2d-1}. A non-Bloch bulk formula, independent of sub-extensive boundary details, can be established only when $\text{dim GBZ}=d$; otherwise, when $\text{dim GBZ}>d$, boundary conditions become crucial for characterizing the system in the thermodynamic limit. We have also implemented this framework on tight-binding lattice Hamiltonians in two and three dimensions and applied it to 2D non-Hermitian reciprocal continuum systems, successfully demonstrating these new phenomena. The outcomes are in complete accord with numerical simulations, exhibiting consistency both quantitatively and qualitatively.
@@
The remainder of the paper is organized as follows. In Sec. II, we summarize key concepts, focusing on dimensional counting to explain the emergence of algebraic skin modes in $d>1$ systems and their transition to exponential decay in cylindrical geometries, without delving into detailed theory analysis. In Sec. III, we introduce rigorous methodologies to derive the GFS framework by employing the standing-wave formation as illustrated in Fig. 1. With this exact theoretical method, we construct the GBZ, compute the open boundary spectrum/eigenstates, and prove the dimension analysis and theoretical framework outlined in Sec. II. In Sec. IV, utilizing this framework, we demonstrate that the non-Hermitian reciprocal systems exhibit the quasi-long-range algebraic skin effect, in which the amplitudes of skin modes follow a power-law decay, instead of the conventionally expected exponential law, when away from the open boundaries. Notably, we emphasize that this algebraic localization is not exclusive to reciprocal skin effects; it can also arise in non-reciprocal skin effects, which are only required to occur in systems with spatial dimension $d\geq 2$. In Sec. V, we show that the algebraic non-Hermitian skin effect is robust even with the presence of random boundary disorders. In Sec. VI, we generalize the GFS formulation into arbitrary $d$ dimensions, and provide examples in three dimensions. In Sec. VII, we extend the above framework from lattice systems to continuum systems.
@@ II Dimensionality analysis for the generalized Brillouin zone and summary of key concepts
Before delving into the universal GFS formula in arbitrary spatial dimensions, we commence with an analysis of dimensionality regarding BZ and GBZ. In this section, our discussion pivots solely on qualitative aspects, leaving the meticulous formulation of the GFS theory to be explored in the subsequent section.
@@
In the construction of eigenstates for a given eigenenergy $E$, one needs to consider only a specific subset of the BZ in Hermitian setups or the GBZ in non-Hermitian systems. This relevant subset will be termed the Fermi surface (FS) in the case of the BZ or the generalized Fermi surface (GFS) in connection with the GBZ. The dimensionality of the FS or GFS can be determined through the following relation:
@@
$\dim{\text{GFS}}$ is a key factor in determining the characteristics of skin modes. When $\dim\text{GFS}=0$, skin modes must decay exponentially with the increase in distance from the open boundaries. On the other hand, if $\dim\text{GFS}>0$, the decay of edge modes may follow different functional forms, such as power-law. Using the dimension counting provided above, it becomes evident that $\dim\text{GFS}>0$ can only occur in dimensions greater than one, $d>1$. It implies that the algebraic non-Hermitian skin effect derived from the nonzero dimension of GFS is exclusive to two and higher dimensional systems. A power-law skin effect at a critical phase transition point in one dimension was reported in Ref. [18], which arises from a distinct mechanism compared to the algebraic skin effect proposed in this work.